# Scalable LLM Math Reasoning Acceleration with Low-rank Distillation

### v1

@@ 4.1 Instruct Models
We first looking into different instruct LLMs which are performative on a variety of non-reasoning tasks and decent reasoning performance. We show Caprese is able to preserve math performance without sacrificing performance on pure language tasks like question answering. In more detail, we test on the Llama 3 family [DJP+24] and Gemma 2 family [TRP+24] of models on zero-shot GSM8K [CKB+21], MATH [HBK+21], CoQA [RCM19], and QASPER [DLB+21], using LM Evaluation Harness [GTA+23] and CoT prompts for math tasks.
Table 1 shows that Caprese is able to preserve most if not all of the math capabilities in the original models without damaging performance on the language tasks, despite the distillation dataset being all math. In most cases, CATS and GRIFFIN severely harm GSM8K and MATH accuracy, but Caprese is able to effectively recover the lost performance. With the exception of Gemma 2 9B Instruct where the layer-wise distillation appears to be better, E2E Caprese is the most performative in the majority of math scenarios, suggesting a benefit beyond minimizing local error. Caprese’s performance is consistent at different sparsity levels and $r$ (Appendix A). The best performers in CoQA and QASPER are a toss-up, but all methods have little impact on the accuracy of these tasks (the main purpose of these tasks is to show no degradation in language-related cases). CATS and Caprese considerably improve GSM8K performance of Llama 3.1 8B Instruct, which has trouble following the prompt template with zero-shot.
@@ 4.2 Scaling Best-of-$N$: Coverage
Figure 3: Coverage and standard deviation of 140 samples from MATH as the number of generation attempts, $N$, scales. We define Relative Compute $\text{Relative Compute Units}=N\times A$ where $A$ is the fraction of total parameters activated per input. FF sparsity set to 50%. Best viewed zoomed.
Now, we see the generalizability of Caprese on the first axis of test-time scaling: sampling multiple responses and selecting the best one, known as Best-of-$N$ (BoN). Increasing the number of generations per prompt, $N$, is one way to scale inference compute for improved performance, as the probability of generating a correct answer increases. To evaluate, we find the coverage (Pass@$K$) on 140 samples from MATH. We use an oracle verifier to accurately assess the quality of the pool of generated responses. To combat the high variance when $K$ is close to $N$, we calculate the average coverage for $K=1,\dots,100$ across 10 independent pools of 100 generations for each sample. Factoring in the saved compute, E2E Caprese is able to have similar or better coverage scaling compared to the original model, as shown in Figure 3. Notably, E2E Caprese (GRIFFIN) on Llama 3.2 3B Instruct improves Pass@$100$ by 7.0% from the original model while using 15.8% less compute.
@@ 4.3 Scaling Length: Thinking Models
Model MATH-500 AIME 2024 AMC 2023
DeepSeek-R1-Distill-Qwen 1.5B 79.40 31.11 $\pm$ 1.92 57.50 $\pm$ 5.00
[…]
Layer-wise Caprese 47.20 3.33 $\pm$ 0.00 27.50 $\pm$ 2.50
E2E Caprese 60.40 6.67 $\pm$ 3.33 35.83 $\pm$ 7.22
[…]
Layer-wise Caprese 73.80 16.67 $\pm$ 3.33 46.67 $\pm$ 2.89
E2E Caprese 74.80 21.11 $\pm$ 1.92 48.33 $\pm$ 3.82
DeepSeek-R1-Distill-Qwen 7B 90.20 51.11 $\pm$ 3.85 76.67 $\pm$ 2.89
[…]
Layer-wise Caprese 84.80 27.78 $\pm$ 1.92 60.83 $\pm$ 1.44
E2E Caprese 85.40 28.89 $\pm$ 6.94 70.00 $\pm$ 5.00
[…]
Layer-wise Caprese 87.00 35.56 $\pm$ 5.09 65.83 $\pm$ 7.22
E2E Caprese 90.00 32.22 $\pm$ 5.09 79.17 $\pm$ 5.77
DeepSeek-R1-Distill-Qwen 14B 92.80 63.33 $\pm$ 3.33 90.83 $\pm$ 2.89
[…]
Layer-wise Caprese 90.80 38.89 $\pm$ 1.92 87.50 $\pm$ 5.00
E2E Caprese 89.20 37.78 $\pm$ 10.72 82.50 $\pm$ 6.61
[…]
Layer-wise Caprese 91.00 60.00 $\pm$ 5.77 88.33 $\pm$ 1.44
E2E Caprese 92.00 58.89 $\pm$ 11.71 85.83 $\pm$ 3.82
Table 2: Thinking models’ 0-shot accuracies on math reasoning tasks. Sparsity is set at 50%, and $r=256$. AIME 2024 and AMC 2023 columns also include the sample standard deviation across 3 runs. Further improvements with reselection are shown in Table 3.
Thinking models provide another axis of test-time scaling by augmenting the CoT length before giving a final answer. Because this entails very long generation lengths, it is critical that error accumulation across tokens be minimized. We test on DeepSeek-R1-Distill-Qwen models [YYZ+24, GYZ+25] using the prompt templates and configurations from Open R1 [Fac25] (temperature and top-$p$ set to 0.6 and 0.95, respectively). For these models, we evaluate zero-shot performance on MATH-500 [HBK+21, LKB+23], AIME 2024, and AMC 2023 for a max generation length of 32768. Due to the small number of problems in AIME 2024 (30 problems) and AMC 2023 (40 problems), we evaluate these 3 independent times, reporting the average accuracy and sample standard deviation.
From Table 2, Caprese is the most performative in most cases. All methods generally appear to be more robust as the model size increases. For instance, CATS and GRIFFIN had very little impact on DeepSeek-R1-Distill-Qwen 14B’s accuracy on MATH-500, with Caprese performing similarly. By far, AIME 2024 is the most challenging and noisy task (due to its difficult and small dataset), with severe degradation from CATS and GRIFFIN and partial recovery using our method. In contrast, AMC 2023 results are less noisy and regained performance from our method is much more substantial. In the next section, we show a simple way to push this recovery even further with reselection.
@@ 4.3.1 Enhanced Performance with Reselection
We can push the performance of Caprese with neuron reselection. For a sample, GRIFFIN and CATS calculate metrics ($\bm{s}$ and $\tau$) to determine subsets of the FF block to use, but these metrics are fixed during the generation phase. After generating many tokens, these selection metrics can benefit from regular updates mid-generation.
@@
For GRIFFIN, updating the metric $\bm{s}$ entails integrating the FF feature statistics of generated tokens into $\bm{s}$. While this can be done by re-running prefill on all tokens, there is a more efficient way. This can be done by passing in the generated tokens following the last reselection through the full model. As these tokens propagate through each layer, we find the selection metric for the generated tokens $\bm{s}_{G}$ and update corresponding KV pairs. This is similar to verification in speculative decoding [CBI+23, LKM23]. Since $\bm{s}$ and $\bm{s}_{G}$ are $\ell_{2}$-norms along the token axis, we can define the updated metric as $\sqrt{(\bm{s}\odot\bm{s})+(\bm{s}_{G}\odot\bm{s}_{G})}$ and use that to reselect different subsets of the FF block to use. This setup updates the pruned layers yet avoids prefill for all tokens. Table 3 shows a clear benefit of reselection (even if infrequent) in Caprese by pushing the performance much closer to the full model.
Model No Reselect $\rho=1024$ $\rho=256$ Full
DeepSeek-R1-Distill-Qwen 1.5B 35.83 $\pm$ 7.22 40.83 $\pm$ 5.77 40.00 $\pm$ 7.50 57.50 $\pm$ 5.00
DeepSeek-R1-Distill-Qwen 7B 70.00 $\pm$ 5.00 68.17 $\pm$ 1.44 74.17 $\pm$ 5.20 76.67 $\pm$ 3.82
DeepSeek-R1-Distill-Qwen 14B 82.50 $\pm$ 6.61 87.50 $\pm$ 4.33 89.17 $\pm$ 3.82 90.83 $\pm$ 2.89
@@
Table 3: E2E Caprese AMC 2023 accuracies and standard deviations when recalculating the GRIFFIN pruning metric during generation. $\rho$ is the rate at which we reselect pruned neurons in terms of generated tokens. No reselection and the full model are special cases where $\rho=\infty$ and $\rho=1$, respectively.
@@
Reselection is also possible with CATS but requires recomputing prefill for all tokens. The parameter to update is the hard thresholding parameter $\tau$, but because this requires finding the desired percentile of intermediate values in the FF block, we would need to access to all values, which likely cannot be fully stored in memory. Consequently, prefill will need to be redone anytime we want to update $\tau$. Additionally, since $\tau$ represents a cutoff for a desired percentile (e.g., median), it is fairly robust to new observations. Given this and the lack of computational incentive, we focus primarily on reselection for GRIFFIN.
4.3.2 Effect on Natural Response Length
Figure 4: Average number of response tokens for MATH-500 queries across different problem difficulties (top) and subjects (bottom). The subjects are algebra, counting & probability, geometry, intermediate algebra, number theory, prealgebra, and precalculus. The global averages are indicated by the dashed lines. Sparsity is set at 50%.
Caprese implicitly encourages brevity, often even producing shorter responses than from the full thinking model. Intriguingly, this behavior arises despite any enforcement or regularization on response lengths anywhere during training or inference (except for cutting off generation at 32K tokens). Shown in Figure 4 with MATH-500, the shortest mean response length for all problem difficulties and subjects is either the full model or Caprese. Meanwhile, CATS consistently outputs the longest responses, averaging roughly 1K more across every sample. It is also interesting to note that with increasing problem difficulty, all models and methods naturally allocate more inference tokens towards answering the question. See Appendix C for similar observations with GRIFFIN.
Given CATS and Caprese (CATS) achieve similar MATH-500 accuracies to the full DeepSeek-R1-Distill-Qwen 7B and 14B models, the ability of Caprese to cut down the lengthy responses of CATS down to the response lengths of the original model or shorter without compromising performance is a direct memory and latency benefit. As we will see in the next section, we can also reduce the per-token latency for further computational savings.
@@ 4.4 Efficiency
Caprese reduces end-to-end generation latency. Table 4 shows the latencies of different setups and models. Caprese cuts latency by >11% when generating 2048 or fewer tokens from a prompt of 2048 tokens using DeepSeek-R1-Distill-Qwen 14B (Qwen 2.5 14B architecture) or Gemma 2 9B. Latency improvements are smaller with generation length 8192 but stilln faster than the full model. Moreover, the latency differences between GRIFFIN and Caprese are fairly small, suggesting that Caprese has minimal overhead. Since we use GRIFFIN on half of each model, the latency reductions are smaller than [DCC24] recorded. Metrics were collected on an NVIDIA L40 GPU using BF16 precision and pre-allocated memory for the KV cache. Time to first/next token metrics are included in Appendix D.
Model Setup Full GRIFFIN Caprese
Qwen 2.5 14B 2048+256 12.8 11.1 ($-13.3\%$) 11.3 ($-11.7\%$)
2048+2048 113.3 98.6 ($-13.0\%$) 100.1 ($-11.7\%$)
2048+8192 661.9 603.2 ($-8.9\%$) 603.7 ($-8.8\%$)
Gemma 2 9B 2048+256 9.6 8.4 ($-12.5\%$) 8.5 ($-11.5\%$)
2048+2048 84.4 74.5 ($-11.7\%$) 75.1 ($-11.0\%$)
2048+8192 483.6 444.4 ($-8.1\%$) 445.8 ($-7.8\%$)
Table 4: End-to-end generation latency (s) for GRIFFIN and Caprese (GRIFFIN). For the “Setup” column, $P+G$ indicates input and generation lengths of $P$ and $G$ tokens, respectively. Percentages in parentheses show the reduction in latency. As before, GRIFFIN is applied to the first half of the model, sparsity is set at 50%, and $r=256$.
@@ Appendix B Caprese Parameters
Appendix C GRIFFIN Response Lengths

### v2

@@ 4.1 Instruct Models
We first look into different instruct LLMs which are performative on non-reasoning tasks and have decent reasoning performance. We show Caprese is able to preserve math performance without sacrificing quality on pure language tasks like question answering. We test on the Llama 3 family [DJP+24] and Gemma 2 family [TRP+24] of models on 0-shot GSM8K [CKB+21], MATH [HBK+21], CoQA [RCM19], and QASPER [DLB+21], using CoT prompts for math tasks.
Table 1 shows that Caprese is able to preserve most if not all of the math capabilities in the original models without damaging performance on the language tasks, despite the distillation dataset being all math. In most cases, CATS and GRIFFIN severely harm GSM8K and MATH accuracy, but Caprese is able to effectively recover the lost performance. E2E Caprese is the most performative in the majority of math scenarios, suggesting a benefit beyond minimizing local error. Caprese’s performance is consistent at different sparsity levels and $r$ (Appendix A). The best performers for CoQA and QASPER are a toss-up, but all methods have little impact on the accuracy of these tasks (the main purpose of these tasks is to show no degradation in language-related cases). CATS and Caprese considerably improve GSM8K performance of Llama 3.1 8B Instruct, which has trouble following the prompt template with zero-shot.
@@ 4.2 Scaling Best-of-$N$: Coverage
Figure 3: Coverage and standard deviation of 140 samples from MATH as the number of generation attempts, NN, scales. We define $\text{Relative Compute Units}=N\times A$ where $A$ is the fraction of total parameters activated per input. $F$ sparsity set to 50%. Best viewed zoomed.
Now, we see the generalizability of Caprese on the first axis of test-time scaling: sampling multiple responses and selecting the best one, known as Best-of-$N$. Increasing the number of generations per prompt, $N$, is one way to scale inference compute for improved performance, as the probability of generating a correct answer increases. To evaluate, we find the coverage (Pass@$K$) on 140 samples from MATH. We use an oracle verifier to accurately assess the quality of the pool of generated responses. To combat the high variance when $K$ is close to $N$, we calculate the average coverage for $K=1,\dots,100$ across 10 independent pools of 100 generations for each sample. Factoring in the saved compute, E2E Caprese is able to have similar or better coverage scaling compared to the original model, as shown in Figure 3. Notably, E2E Caprese (GRIFFIN) on Llama 3.2 3B Instruct improves Pass@$100$ by 7.0% from the original model while using 15.8% less compute.
@@ 4.3 Scaling Length: Thinking Models
Model MATH-500 AIME 2024 AMC 2023 BRUMO 2025
DS-Qwen 1.5B 79.40 30.00 $\pm$ 2.72 60.83 $\pm$ 7.47 25.83 $\pm$ 3.19
GRIFFIN 42.00 2.50 $\pm$ 1.67 16.88 $\pm$ 1.25 1.67 $\pm$ 1.92
Layer-wise Caprese 47.20 2.50 $\pm$ 1.67 28.13 $\pm$ 2.39 0.00 $\pm$ 0.00
E2E Caprese 60.40 6.67 $\pm$ 2.72 34.38 $\pm$ 6.57 4.17 $\pm$ 1.67
CATS 72.00 11.67 $\pm$ 6.94 37.50 $\pm$ 3.54 10.83 $\pm$ 4.19
Layer-wise Caprese 73.80 15.83 $\pm$ 3.19 48.13 $\pm$ 3.75 13.33 $\pm$ 2.72
E2E Caprese 74.80 20.83 $\pm$ 1.67 47.50 $\pm$ 3.54 20.00 $\pm$ 2.72
DS-Qwen 7B 90.20 47.50 $\pm$ 7.87 75.00 $\pm$ 4.08 45.83 $\pm$ 4.19
GRIFFIN 80.60 21.67 $\pm$ 4.30 62.50 $\pm$ 8.90 25.00 $\pm$ 8.39
Layer-wise Caprese 84.80 29.17 $\pm$ 3.19 64.38 $\pm$ 7.18 27.50 $\pm$ 4.19
E2E Caprese 85.40 30.00 $\pm$ 6.09 71.88 $\pm$ 5.54 29.17 $\pm$ 5.69
CATS 89.20 34.17 $\pm$ 7.88 62.50 $\pm$ 2.04 37.50 $\pm$ 7.39
Layer-wise Caprese 87.00 35.00 $\pm$ 4.30 70.00 $\pm$ 10.21 39.17 $\pm$ 3.19
E2E Caprese 90.00 33.33 $\pm$ 4.71 78.13 $\pm$ 5.15 47.50 $\pm$ 5.00
DS-Qwen 14B 92.80 63.33 $\pm$ 2.72 90.63 $\pm$ 2.39 60.83 $\pm$ 4.19
GRIFFIN 89.80 32.50 $\pm$ 7.39 80.63 $\pm$ 3.75 33.33 $\pm$ 3.85
Layer-wise Caprese 90.80 41.67 $\pm$ 5.77 86.88 $\pm$ 4.27 43.33 $\pm$ 4.71
E2E Caprese 89.20 39.17 $\pm$ 9.18 84.38 $\pm$ 6.57 40.00 $\pm$ 2.72
CATS 92.80 58.34 $\pm$ 1.92 88.75 $\pm$ 1.44 55.00 $\pm$ 3.33
Layer-wise Caprese 91.00 61.67 $\pm$ 5.77 88.75 $\pm$ 1.44 53.00 $\pm$ 7.20
E2E Caprese 92.00 58.34 $\pm$ 9.62 87.50 $\pm$ 2.04 49.17 $\pm$ 4.19
Table 2: Thinking models’ 0-shot accuracies on math reasoning tasks. Sparsity is set at 50%, and $r=256$. AIME 2024 and AMC 2023 columns also include the sample standard deviation across 4 runs. Further improvements with reselection are shown in Table 3.
Thinking models provide another axis of test-time scaling by augmenting the CoT length before giving a final answer. Because this entails very long generation lengths, it is critical that error accumulation across tokens be minimized. We test on DeepSeek-R1-Distill-Qwen (which we abbreviate to DS-Qwen) models [YYZ+24, GYZ+25] using the prompt templates and configurations from Open R1 [Fac25] (temperature and top-$p$ set to 0.6 and 0.95, respectively). For these models, we test 0-shot performance on MATH-500 [HBK+21, LKB+23], AIME 2024 [AIM25], AMC 2023 [AMC23], and BRUMO 2025 [BRU25] for a max generation length of 32768. Due to the small size of AIME 2024, AMC 2023, and BRUMO 2025, we evaluate them 4 times, reporting the average accuracy and sample standard deviation.
From Table 2, Caprese is the most performative in most cases. DS-Qwen 1.5B and 7B see the greatest benefit with E2E Caprese usually achieving the highest accuracy, sometimes even exceeding the full model’s performance. DS-Qwen 7B with CATS brings AMC 2023 accuracy down to 62.50% from 75.00%, but Caprese lifts performance above the original model’s to 78.13%, and similarly with DS-Qwen 7B with CATS and BRUMO 2025. All methods are more robust as the model size increases. For instance, CATS and GRIFFIN had very little impact on DS-Qwen 14B’s accuracy on MATH-500, with Caprese performing similarly. AIME 2024 and BRUMO 2025 challenge CATS and GRIFFIN with several degradation and partial recovery using our method. In the next section, we show a simple way to push this recovery even further with reselection.
@@ 4.3.1 Enhanced Performance with Reselection
We can push the performance of Caprese with neuron reselection. For a sample, GRIFFIN and CATS calculate metrics ($\bm{s}$ and $\tau$) to determine subsets of the FF block to use, but these metrics are fixed during generation. Updating them mid-generation can benefit downstream performance.
@@
For GRIFFIN, updating the metric $\bm{s}$ entails integrating the FF feature statistics of generated tokens into $\bm{s}$. While this can be done by re-running prefill on all tokens, there is a more efficient way by passing in the generated tokens following the last reselection through the full model. As these tokens propagate through each layer, we find the selection metric for the generated tokens $\bm{s}_{G}$ and update corresponding KV pairs. This is like verification in speculative decoding [CBI+23, LKM23]. As $\bm{s}$ and $\bm{s}_{G}$ are $\ell_{2}$-norms along the token axis, we define the updated metric as $\sqrt{(\bm{s}\odot\bm{s})+(\bm{s}_{G}\odot\bm{s}_{G})}$ and use that to reselect different subsets of the FF block to use (as described in Section 2). This updates the pruned layers yet avoids prefill for all tokens. Table 3 shows a clear benefit of reselection (even if infrequent compared to speculative decoding) in Caprese by pushing the performance much closer to the full model’s as we decrease the number of decode steps between reselection rounds. Although GRIFFIN is generally more harmful to accuracy than CATS due to its structured pruning, paired with Caprese and reselection, Caprese (GRIFFIN) can exceed the performance of Caprese (CATS) and reach the full model performance.
Model No Reselect $\rho=1024$ $\rho=256$ $\rho=64$ Full
DS-Qwen 1.5B 34.38 41.87 45.00 58.75 60.83
DS-Qwen 7B 71.88 69.25 73.13 73.75 75.00
DS-Qwen 14B 84.38 88.13 88.75 91.88 90.63
@@
Table 3: E2E Caprese AMC 2023 accuracies and standard deviations when recalculating GRIFFIN pruning metrics and reselecting pruned neurons every $\rho$ decode steps. No reselection and the full model are special cases where $\rho=\infty$ and $\rho=0$, respectively.
@@
Reselection is also possible with CATS but requires recomputing prefill for all tokens. The parameter to update is the hard thresholding parameter $\tau$, but because this requires finding the desired percentile of intermediate values in the FF block, we would need to access to all values, which likely cannot be fully stored in memory. In turn, prefill will need to be redone anytime we want to update $\tau$. Since $\tau$ represents a cutoff for a desired percentile (e.g., median), it is fairly robust to new observations. Given this and the lack of computational incentive, we focus primarily on reselection for GRIFFIN.
@@ 4.4 Efficiency
Caprese reduces generation latency. Table 4 shows the latencies of different generation settings. Caprese cuts total latency and time to next token by >16% for the Qwen 2.5 14B architecture. Moreover, the latency differences between GRIFFIN and Caprese are exceptionally small, suggesting that Caprese has minimal overhead. We also include latencies of GRIFFIN paired with LoRA weights, which although faster than the base model, negates >40% of the time savings that GRIFFIN has provided. This highlights the efficiency suboptimality of LoRA during inference. Metrics were collected on an NVIDIA L40 GPU using BF16 precision.
Setup End-to-end Latency (s) Avg. Time to Next Token (ms)
Full GRIFFIN LoRA Caprese Full GRIFFIN LoRA Caprese
2048+256 10.5 8.7 9.5 8.7 41 34 37 34
2048+2048 84.4 70.3 76.5 70.4 41 34 37 34
2048+8192 344.9 287.7 312.5 288.8 42 35 38 35
Table 4: End-to-end generation latency (s) and average time to next token (ms) for GRIFFIN, LoRA, and Caprese using Qwen 2.5 14B. For the “Setup” column, $P+G$ indicates input and generation lengths of $P$ and $G$ tokens, respectively. As before, GRIFFIN is applied to the first half of the model, sparsity is 50%, and $r=256$.
4.4.1 Bonus: Effect on Natural Response Length
Figure 4: Average number of response tokens for MATH-500 queries with increasing problem difficulty. The global averages are indicated by the dashed lines. Sparsity is set at 50%.
Caprese implicitly encourages brevity, often even producing shorter responses than from the full thinking model. Intriguingly, this behavior arises despite any enforcement or regularization on response lengths anywhere during training or inference (except for cutting off generation at 32K tokens). Shown in Figure 4 with MATH-500, the shortest mean response length for all problem difficulties is either the full model or Caprese. Meanwhile, CATS consistently outputs the longest responses, averaging roughly 1K more across every sample. It is also interesting to note that with increasing problem difficulty, all models and methods naturally allocate more inference tokens towards answering the question. Given CATS and Caprese (CATS) achieve similar MATH-500 accuracies to the full DS-Qwen 7B and 14B models, the ability of Caprese to cut down the lengthy responses of CATS down to the response lengths of the original model or shorter without compromising performance is a direct memory and latency benefit. This brevity of Caprese is more pronounced in larger models, proportionally, from a 5.8% reduction in tokens compared to the full DS-Qwen 7B model to an 8.5% reduction compared to the full DS-Qwen 14B model. The cause of this phenomenon is of interest for future work. See Appendix D for similar observations with GRIFFIN.
@@ Appendix B Caprese Parameters
Appendix C LoRA Performance
We report the performance of using LoRA in place of Caprese while keeping parameter count constant in Table 5. Caprese achieves the highest accuracy in slightly more cases than LoRA, but as per our discussion on LoRA in Section 3.5 and efficiency analysis in Section 4.4, LoRA’s main benefit is efficient training, not efficient inference.
Model GSM8K MATH CoQA QASPER
Llama 3.2 1B Instruct 22.44 10.66 55.43 14.43
GRIFFIN 7.13 5.42 56.05 14.11
Layer-wise LoRA 8.11 5.60 56.00 14.65
E2E LoRA 20.47 10.10 56.10 14.41
Layer-wise Caprese 13.72 6.62 56.07 13.40
E2E Caprese 21.00 8.44 56.55 13.88
Llama 3.2 3B Instruct 51.55 14.32 63.95 12.45
GRIFFIN 28.96 10.98 64.52 12.52
Layer-wise LoRA 26.46 11.28 63.88 12.55
E2E LoRA 42.91 16.50 64.67 12.63
Layer-wise Caprese 40.18 13.70 64.33 11.60
E2E Caprese 44.66 16.96 64.83 12.35
Gemma 2 2B Instruct 51.02 16.06 63.77 10.96
GRIFFIN 33.74 11.32 63.28 11.07
Layer-wise LoRA 39.58 14.66 62.98 10.54
E2E LoRA 44.66 16.71 63.03 10.81
Layer-wise Caprese 42.53 12.32 63.77 10.75
E2E Caprese 48.14 13.70 63.37 11.05
Table 5: 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA and QASPER) with LoRA. $F$ sparsity is set to 50%, $r=256$, and LoRA ranks are set to match the number of parameters that Caprese adds.
Appendix D GRIFFIN Response Lengths